2019-04-12-LeadershipReports

From Administrative
Jump to: navigation, search

April 12, 2019 Report - April 1 - April 26 Director - Todd Satogata

  Previous two weeks (Apr 1 - Apr 12)
     - 
  Next two weeks (Apr 15 - Apr 26)
     - 

Alex Bogacz

  Previous two weeks (Apr 1 - Apr 12)
     - Machine support 
     - Paper on ‘Dogbone’ ERL for IPAC
     - Finalizing White Paper on Publications and Conferences
     - JLEIC Collab. mgt.
  Next two weeks (Apr 15 - Apr 26)
     - Injector solenoids - importance of uniformity of focusing across the bunch and stray field 
     - Finalizing paper on ‘Dogbone’ ERL for IPAC
     - Implementation of new Internal Publication system
     - PRAB paper - review

Rui Li

  Previous two weeks (Apr 1 - Apr 12)
     - Participated in the EIC collaboration meeting, presented my talk, and interacted with colleagues 
       from different labs
     - Extended discussions with the RF group and John Fox and Dmitry Teytelman on feedback requirements for JLEIC
     - Scheduling the next impedance meeting and update impedance wiki 
     - Started collaboration with LBL on impedance and wakefield studies for ALS-U
     - Literature studies on impedances
  Next two weeks (Apr 15 - Apr 26)
     - Continue with broadband impedance study for JLEIC
     - Impedance meeting on 4/22 at 10:30am
     - Review some of the studies on CSR effects 

Ed Nissen

  Previous two weeks (Apr 1 - Apr 12)
     - Attended the JLEIC collaboration meeting 
     - Working on ECA application
  Next two weeks (Apr 15 - Apr 26)
     - Continue working on ECA application
     - Finish paper to submit for light peer review at IPAC

Chris Tennant

  Previous two weeks (Apr 1 - Apr 12)
     - JLEIC collaboration meeting
     - Finished collecting parameters for updated ERL landscape plot (with help of Peter Williams)
     - Met with SLAC representatives about collaborating on their LDRD (ML relating to SRF)
     - Went over SRF classification analysis with Anna and Adam
     - Continue analysis and documentation in TN for SRF classification (text needed for Anna's LDRD, 
       IPAC paper and Adam's ICALEPCS paper)
     - Vacation day to attend "Applied Machine Learning Conference" in Charlottesville
  Next two weeks (Apr 15 - Apr 26)
     - Finish(?) TN



Accelerator R & D - Yuhong Zhang

Fanglei Lin

  Previous two weeks (Apr 1 - Apr 12)
     - Maternity Leave
     - Finished the electron collider ring design
     - Updating the pCDR-100 in the electron complex section
     - Prepared electron ring information to the talk given in the collaboration meeting spring 2019 
     - Provided the electron ring magnet information to Leigh and Tim for the cost estimation
     - Had a discussion and made a plan on the study of electron polarization with colleagues from UNM 
       and Cornell University 
  Next two weeks (Apr 15 - Apr 26)
     - Maternity Leave


Vasiliy Morozov

  Previous two weeks (Apr 1 - Apr 12)
     - JLEIC collaboration meeting: preparation and presentation of four talks on JLEIC overview, IR design, collider
       ring designs and working group summary, discussions with collaborators. Presentation of the latest JLEIC 
       design to NP.
     - Costing and energy upgrade exercises.
     - Ion polarization: spin transparency experiment analysis and simulations, preparation of a paper with Kondratenko.
     - IR design: setting up collider ring simulations in Elegant with Randy.
     - Work on simulations of the longitudinal dynamics in a storage ring cooler with Bhawin.
     - Work on beam-beam simulations with River
  Next two weeks (Apr 15 - Apr 26)
     - Costing exercise.
     - Preparation of the DoE quarterly reports.
     - Work on the spin transparency experiment and polarization papers
     - Completion of the pCDR.
     - Collider ring design and simulations.
 


Amy Sy

  Previous two weeks (Apr 1 - Apr 12)
     - Attended JLEIC collaboration meeting.
     - Continued work on early career proposal.
     - Participated in commissioning of laser-driven polarized proton source at FZ Juelich. Testing Lamb-shift
       polarimeter
       with unpolarized pulsed proton beam.
     - Summarized longitudinal Schottky signal analysis work on IMP experiment data for quarterly report.
  Next two weeks (Apr 15 - Apr 26)
     - Continued work on early career proposal.
     - Continued participation in experiments with Lamb-shift polarimeters at FZ Juelich.


Yuhong Zhang

  Previous two weeks (Apr 1 - Apr 12)
     - Organized and attended JLEIC collaboration meeting.
     - Met with various visitors during the week of the collaboration meeting
     - Activites supporting the cost exercises (design and risk factors)
     - Worked on various parts of the JLEIC design, particularly 140 GeV high energy JLEIC
  Next two weeks (Apr 15 - Apr 26)
     - Continue work on various parts of the JLEIC design, particularly 140 GeV high energy JLEIC



Computational Physics - Yves Roblin
Yves Roblin

   CEBAF
     - Prepare for machine startup.
     - Preparing talk to Hall A collaboration.
     - Completed TOSP for Hall A Apex Experiment
     - Developing method to center solenoids (in particular 
        Hall A moeller solenoid)
     - BTEAM
    JLEIC:   
     - Reviewing Implications of the 100 GeV CM setup for BeamBeam   
   OTHER
     - Attended workshop on ML
     - Conducted interviews for postdoc position
     - TRC meetings
     - FOA reports.
     - Evaluating various Python frameworks for ML and GA (DEAP,        pyEVOLVE)


He Zhang

   Previous two weeks (Apr 1 - Apr 12)
     - MAD-X 5D Map tracking code has been sent to Yun Luo and Ji Qiang
     - Working on JSPEC development
   Next two weeks (Apr 15 - Apr 26)
     - Enhance the luminosity module of JSPEC
     - Simulation on dispersive cooling


River Huang

   Previous two weeks (Apr 1 - Apr 12)
     - “CASA Beam-Beam” update: Updated the interface between ““CASA Beam-Beam” and “Elegant””: the interface-script 
        file under Linux system was translated by python and the python code was added to “CASA Beam-Beam”. Now 
        “CASA Beam-Beam” can generate “mtr” and “lte” format-file automatically; “CASA Beam-Beam” can automatically
        generate “ele” files of Elegant command, again, the results of “CASA Beam-Beam” can be used for beam-beam 
        interaction in Elegant as well. Optimized the demonstration windows of “CASA Beam-Beam” GUI so that the user
        can review numerical results; added the script commands into “CASA Beam-Beam” to plot/save the figure of “Spin
        Response Function”.
      - Simulating and studying - We are simulating and studying the effect on bunch length due to synchrotron tune via 
        both CASA beam-Beam and BeamBeam3D.
      - Attended “JLEIC collaboration meeting,” reported on "Update on JLEIC Ion Polarization and Spin 
        Transparent Experiment"
      - Reported on “Update on JLEIC Ion Polarization and Spin Transparent Experiment”.
   Next two weeks (Apr 15 - Apr 26)
      - Continue to simulate of the zero-integer spin resonance strength and analyze the results;
      - Study and simulate “Beam-Beam effect”;
      - Prepare the poster for IPAC19.



Diagnostic Development - Kevin Jordan

   Previous two weeks (Apr 1 - Apr 12)
     - Attended an ARIES sponsored workshop on scintillating screens, cameras & optics.
     - Working on getting power setup for Magnetron Project. This will be done in Lab 1 of the LERF. Within the next
       few weeks the first of two transmitters will arrive and be installed in the LERF.
     - Mounted a BNNT Halo monitor screen for the Isotope beam line run the LERF, also worked to get beam to 0G dump 
       and to the 1G dump. Initially the vacuum system was not ready then we were all ready & the drive laser would 
       not stay locked. There were a num bet of items on HCO that were signed off but not operational. We will try to 
       run to the 1G dump in ~6 weeks when we are back at 2K.
     - Working with Joe & Mike to get the first SLM installed in Arc 7 during the SAD. I also mounted a BNNT flag as 
       a part of this installation. 
     - Attending a meeting at ODU about hope to better work together on common projects with the school of engineering.
     - Finishing up some IBIC program details for this fall’s conference.
     - Spent some time on the IPM gun pulser design for the 300 kV experiment.
   Next two weeks (Apr 15 - Apr 26)
     - 


Joe Gubeli

   Previous two weeks (Apr 1 - Apr 12)
     - Finished and implemented the new optical layout of ARC 7 SLM. We now have a filter wheel in front of the first
       optic as Mike T. asked for the second filter wheel is 90mm upstream of the third filter wheel which is located at
       the field stop. The first wheel will contain the slits. I am waiting on input from Mike T. on the width and 
       offset of the five different slits that will be installed here. The second wheel will contain bandpass filters 
       for short and long visible wavelengths along with a vertical and a horizontal polarizer. The last wheel contains
       neutral density filters. Met with Steve S. and Keith C. at the installation location. Steve said he will approve
       the installation but hasn’t yet updated the Atlis.
     - Spent time working on the Beamenizer after we found the system not working when running to the 0G dump. I created
       a test fixture and after a timing board repair I was able to verify the system now works.
   Next two weeks (Apr 15 - Apr 26)
     - VACATION - 1 week
     - Prepare and install the ARC 7 SLM.


Michael Tiefenback

   Previous two weeks (Apr 1 - Apr 12)
     - Specify quadrupole configuration for Hall B for summer run (HPS beam spots ~ 30 microns)
     - Specify depth of field limiting slit widths for Arc 7 SLM
     - Specify dual-magnet steering modulation for real-time rayTrace (in progress; appears promising)
     - Planning for overall diagnostic system integration, including periodic calibration requirements
   Next two weeks (Apr 15 - Apr 26)
     - model SyncLightMonitor matching process for arcs
     - Review field map updates for Moller quads requested by Benesch
     - Continue on 4) above (diagnostic system integration)



LERF - Steve Benson

   Previous two weeks (Apr 1 - Apr 12)
     - Attended the Spring JLEIC Collaboration meeting and hosted the cooling working group. 
     - Presented Cooler ring status presentation at the collaboration meeting. 
     - Completed run plan for running beam to the 1G dump.
     - Tried to run beam to the 1G dump but the drive laser failed so no beam was run. 
     - Hosted Cooling ring  and FOA status meeting and wrote up minutes.
     - Rewrote and submitted the LOSP for the drive laser alignment in the LERF vault.
   Next two weeks (Apr 15 - Apr 26)
     - Complete the quarterly report for the FOA Cooling project.
     - Complete the budget for the Isotope project quarterly.
     - Submit new text for LOD and safety envelope for LERF.
     - Summarize the ERLs for EIC workshop.